To warm up for a match, a tennis player hits the 58.0 g ball vertically with her racket.

If the ball is stationary just before it is hit and goes 5.10 m high, what impulse did she impart to it?
Answer #1

According to the question we need to find the impulse we know the formula for impulse .

Impulse = mass * ( v(i) )

We use the equation

V(f)^2 = V(i) ^2 + 2as

V(i) = sqrt ( 2 * 9.8 * 5.10 ) = 10.0 m/s

Plug value in impulse formula

Impulse = 0.058 kg * 10.0 m/s = 0.58 kg m/s ========================ANSWER)